Circular No. 8975

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 R6 (GIACOBINI) S. Nakano, Sumoto, Japan, reports the discovery of a comet by Koichi

Itagaki (Teppo-cho, Yamagata) and Hiroshi Kaneda (Minami-ku, Sapporo) on unfiltered CCD patrol frames taken on Sept. 10.56 UT using a 0.21-m f/3 reflector; the position tabulated below was measured from a confirming unfiltered CCD image taken subsequently with a 0.60-m f/5.7 reflector at Yamagata, noting that the comet is diffuse with strong condensation, with a coma diameter of ∼ 25′′ and a 2′ tail toward the east-southeast.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.10.56854  20^{h}39^{m}18.^{s}70  −1^{o}17′39′′.9  13.5

Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have noted the cometary appearance of this object (some of the reports being sent by Nakano), with the coma diameter generally given as 20′′-50′′ and the tail length given as 20′′-66′′ in p.a. ∼ 106^{o}–124^{o} (such reports have been received from Y. Ikari, Moriyama, Shiga-ken, Japan; K. Kadota, Ageo, Saitama-ken, Japan; V. Gonano and E. Guido, Remanzacco, Italy; R. Ligustri, Talmassons, Udine, Italy; J. J. Gómez, Gandia, Spain; J. A. Henríquez, Tenerife, Spain; P. Camilleri and G. Sostero, remotely using a 0.30-m reflector at Catania, Italy; and F. Castellani, Monte Baldo, Italy).

M. Meyer, Limburg, Germany, suggests that this comet is identical to comet D/1896 R2 (Giacobini), which has not been seen since Jan. 1897 and for which a prediction by Nakano (2005, Nakano Note 1211, based on an orbit by Belyaev et al. 1974, QJRAS 15, 450) gives T = 2008 Sept. 9.89 TT. The identity has been confirmed by Nakano, who notes that the comet has made 17 revolutions and passed only 0.51 AU from the earth on 1962 Sept. 9 and 0.81 AU from Jupiter on 1992 Jan. 14. The following linked orbital elements by B. G. Marsden are from 93 observations, 1896–2008 (including newly reduced astrometry by G. V. Williams from 1896–1897); these elements, together with the 1896 elements and an ephemeris, appear on MPEC 2008-R60.

Epoch = 2008 Sept. 11.0 TT

T = 2008 Sept.10.2063 TT  ω = 154.^{o}3051
e = 0.568403  Ω = 179.6275 }2000.0
q = 1.527870 AU  i = 15.3177

a = 3.540037 AU n^{o} = 0.1479764 P = 6.66 years

2008 September 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green